OpenStudy (anonymous):

4x-2y=-2, 6x+y=5 solve by using the elimination method

OpenStudy (anonymous):

4x-2y=-2, 6x+y=5 solve by using the elimination method

OpenStudy (tasreeb):

4x-2y=-2 6x+y=5 multiply by 2 eq:2 12x+2y=10 16x=8 x=8/16 x=1/2

OpenStudy (anonymous):

 @tasreeb has got that correct :) now plug into the other equation